OpenStudy (candeii):

Omg someone help in desperate need for help I don't understand radicals. So my teacher said we have to find the greatest factor in order to get the problem right? Ex: rad 20 / Factors of 20: 4•5 (We found 4 as a perfect square root) / rad 4 • rad 5 = 2 rad five. Now these are the ones I need help on! 1. Rad 1100 2. Rad 63 3. 3 rad 121 Thank you !!

jhonyy9 (jhonyy9):

1. 11*100 = 11*10^2 = 10 sqrt 11 2. rad63 = rad(9*7) = rad(3^2 *79 = 3rad 7 3. 3rad121 = 3rad(11*11) = 3rad(11^2) = 3*11 = 33 hope helped you and this is understandably sure easy